OpenStudy (anonymous):

FAN AND MEDAL Amir pitches a baseball at an initial height of 6 feet, with a velocity of 73 feet per second. If the batter misses, about how long does it take the ball hit the ground? Hint: Use H(t) = −16t^2 + vt + s. 4.64 seconds 2.94 seconds 2.28 seconds 0.08 second

OpenStudy (anonymous):

@queen-of-tokyo @william22 @enchanted_bubbles @rational @triciaal @yomamabf @uybuyvf @ikram002p @oldrin.bataku @poopsiedoodle @Astrophysics @satellite73 @DullJackel09 @freckles @ganeshie8 @hartnn @Hero @JFraser @Kainui @kropot72 @Luigi0210 @zepdrix @campbell_st @vera_ewing

OpenStudy (anonymous):

@hartnn help please!

hartnn (hartnn):

just plug in the values given in the equation, H(t) = −16t^2 + vt + s. here s= initial height = 6 v = velocity = 73 and the height at ground level H(t) = 0 only 't' is unknown!

hartnn (hartnn):

you'll get a quadratic in 't' let me know if you have any doubts in solving that?

OpenStudy (anonymous):

how do i find t?

hartnn (hartnn):

did u plug in values? what did u get?

OpenStudy (anonymous):

-16t^2+73t+6=0

hartnn (hartnn):

now use quadratic formula to solve that!

hartnn (hartnn):

Compare your quadratic equation with \(ax^2+bx+c=0\) find \[a=...?\\b=...?\\c=...?\\\] \[ \\ \sqrt{b^2-4ac}=...?\] then the two roots of x are: \(\huge{x_{1,2}=\frac{-b \pm \sqrt{b^2-4ac}}{2a}}\)

OpenStudy (anonymous):

oh ok can you wait for me to solve it?

OpenStudy (triciaal):

don't tag me anymore

hartnn (hartnn):

sure, just gave you the steps :)

OpenStudy (anonymous):

y? @triciaal

OpenStudy (anonymous):

alright I'm almost done

OpenStudy (anonymous):

is it 4.64? @hartnn

hartnn (hartnn):

yes! correct :)

OpenStudy (anonymous):

thank you!!

hartnn (hartnn):

welcome ^_^